# Quantum synchronization of twin limit-cycle oscillators

### v1

@@ Abstract
Introduction—Synchronization phenomena of self-sustained oscillating systems have been studied for centuries [1, 2] and remain an active research topic of non-linear dynamics [3, 4, 5, 6]. Self-sustained oscillators exhibit limit cycles, i.e., closed phase-space trajectories that the system converges to regardless of its initial state. Such limit-cycle oscillators can align quantities like their phase or frequency with other oscillators as well as with external signals. Quantum analogues of limit-cycle oscillators have been studied in various settings ranging from quantum van der Pol oscillators [7, 8, 9, 10, 11, 12] to few-level quantum oscillators [13, 14, 15, 16]. While some features of quantum synchronization are inherited from classical synchronization, others originate from genuine quantum properties like entanglement [13, 17, 18, 19, 20, 21] and (destructive) interference that can lead to a synchronization blockade [22, 23, 24]. Understanding quantum synchronization has implications for quantum sensing [25], quantum thermodynamics [26, 27, 28], quantum networks [29], and time crystals [30, 31, 32]. First steps towards its experimental realization have been taken on several platforms including cold atoms [33], nuclear spins [34], trapped ions [35], and superconducting circuits [36].
@@
While there has been an increasing interest in quantum systems featuring multiple separate limit cycles [40, 41, 42, 43, 44], their synchronization properties have not been studied. In this work, we introduce a Liénard system where two limit cycles coexist in a single steady state, regardless of the initial conditions and investigate their synchronization behavior. We call this system a twin limit cycle (TLC): it is characterized by a double ring-like structure in phase space as sketched in Fig. 1(b). The location of the minima (maxima) of the effective potential of the corresponding classical Liénard system is indicated by the dashed (dotted) rings. Our setup can be extended to host multiple limit cycles. We examine the synchronization of a single TLC under an external coherent drive and find that the limit cycles exhibit different locking behaviors. Furthermore, for two coupled identical TLCs, both synchronization and blockade effects coexist, an apparent paradoxical interplay unattainable with standard limit cycles. To distinguish the contributions of individual limit cycles of a TLC, we define new finer measures of quantum synchronization. Finally, we outline an experimental setup to implement our model.
Figure 1: Illustration of a system with multiple limit cycles. (a) Sketch of the effective potential $V(r)$ of a classical Liénard system featuring two basins of attraction separated at $r_{c}$. (b) Wigner function of a twin limit cycle/quantum Liénard system with $\langle a\rangle=r\mathrm{e}^{\mathrm{i}\phi}$. The dashed (dotted) rings with radii $r_{1}$ and $r_{2}$ ($r_{c}$) correspond to the local minima (maximum) of the effective potential obtained from the mean-field equations of motion, see Eq. 4.
Model— We consider a coherently driven anharmonic quantum oscillator subject to incoherent first and third-order pumping, along with second and fourth-order damping. These dissipative processes stabilize two concentric limit cycles, see Fig. 1(b). The dynamics in the rotated frame of the drive is described by the master equation
@@
where $H_{0}=\Delta a^{\dagger}a^{\phantom{\dagger}}+Ka^{\dagger 2}a^{2}$, $H_{d}=\Omega(a+a^{\dagger})$, and $a$ ($a^{\dagger}$) denote the annihilation (creation) operators of the oscillator. The detuning between the TLC and the drive is denoted by $\Delta$, $K$ parametrizes the Kerr nonlinearity, and $\Omega$ denotes the strength of the drive. The rates $\gamma_{j}$ correspond to incoherent gain (odd $j$) and damping (even $j$). The model simplifies to the paradigmatic quantum van der Pol oscillator with a single limit cycle for $\gamma_{3}=\gamma_{4}=0$. Additional incoherent processes of higher order, $\mathcal{D}[a^{\dagger n}]\rho$ and $\mathcal{D}[a^{m}]\rho$, with $n(m)$ being odd(even), can be included to obtain multiple limit cycles.
We begin by examining the semiclassical limit to obtain an approximation to the steady state of the quantum system. The mean-field equations of the effective classical Liénard model can be derived by setting $\langle a\rangle=r\mathrm{e}^{\mathrm{i}\phi}$, resulting in
@@
ϕ˙=˙italic-$\dot{\phi}=$ $-\Delta-2Kr^{2}-\frac{\Omega}{r}\cos(\phi)\,.$ (3)
For a vanishing drive $\Omega=0$ and suitable choice of the rates $\gamma_{j}$, Eq. 2 can be expressed in terms of its three roots $r_{1}<r_{c}<r_{2}$,
$\dot{r}=$ $r(r_{1}^{2}-r^{2})(r_{c}^{2}-r^{2})(r_{2}^{2}-r^{2})2\gamma_{4}\equiv-\partial_{r}V(r)\,,$ (4)
where $r_{1}$ and $r_{2}$ are the stable solutions, $r_{c}$ is the unstable solution of the mean-field equations, and $V(r)$ is the effective potential, see Fig. 1(a). In a classical system, $r_{c}$ separates the two basins of attraction. An initial state with $r<r_{c}$ ($r>r_{c}$) will therefore converge to $r_{1}$ ($r_{2}$).
@@
In a classical Liénard system of two limit cycles, both basins of attraction are separated at $r_{c}$. However, if extrinsic noise is added, trajectories can cross this boundary. In the language of the effective potential shown in Fig. 1(a), noise-induced jumps in $r$ have to overcome the potential barriers $\Delta V_{j}$ to “tunnel” between both limit cycles. In our quantum setup, the steady state is a combination of two distinct quantum limit cycles. Considering quantum trajectories, the system tunnels between the two limit cycles due to inherent quantum noise, see [45] for a detailed discussion.
@@
Figure 2: Drive-induced phase locking of a twin limit cycle (TLC). (a) Wigner function for $\Omega=8\gamma_{1}$ and $\Delta=K=0$ showing phase localization of both limit cycles, i.e., a maximum at $\phi=\mathrm{arg}(\langle a\rangle)=-\pi/2$. Here, dashed (dotted) rings correspond to stable (unstable) solutions of the mean-field equations of the undriven limit cycles, see Eq. 4. Note the power-law color scale. In panels (b) and (c), solid curves denote the maximum of $P_{1}^{\alpha}$ with $\alpha\in\{\mathrm{in,out}\}$, dashed curves denote $\text{arg}(\langle\tilde{a}_{\alpha}\rangle)$. Varying $\Delta$ and $K$, the inner limit cycle exhibits a larger phase shift than the outer one. This is opposite to the behavior of a standard quantum van der Pol oscillator. Here, $\Omega=0.25\gamma_{1}$ for both (b) $K=0$ and (c) $\Delta=0$. The dissipation rates for (a)–(c) are $\gamma_{2}=2.5\gamma_{1}$, $\gamma_{3}=1.04\gamma_{1}$, and $\gamma_{4}=0.096\gamma_{1}$.
One driven twin limit cycle— We first focus on the synchronization properties of this quantum Liénard system. First, we discuss phase locking of a TLC to an external drive. In Fig. 2(a), we present the Wigner function corresponding to the steady state of a driven TLC, which exhibits phase localization near $\phi=\mathrm{arg}(\langle a\rangle)=-\pi/2$ indicating the synchronization of both limit cycles to the external drive. To characterize and quantify the amount of synchronization, we define a phase localization measure. Various measures of quantum synchronization have been proposed in the literature [7, 13, 46, 47, 48]. In this work, we make use of the synchronization measure based on phase states |$\ket{\phi}=\sum_{n=0}^{\infty}\mathrm{e}^{\mathrm{i}n\phi}\ket{n}/\sqrt{2\pi}$ [49, 47] defined as
$P_{1}(\phi)=\bra{\phi}\rho\ket{\phi}-\frac{1}{2\pi}=\frac{1}{2\pi}\sum_{n,m=0}^{\infty}\mathrm{e}^{\mathrm{i}(m-n)\phi}\rho_{nm}-\frac{1}{2\pi}\,,$ (5)
where $\rho_{nm}=\bra{n}\rho\ket{m}$. This measure can be further simplified to
$P_{1}(\phi)=$ $=\frac{1}{2\pi}\sum_{k=1}^{\infty}\mathrm{e}^{-\mathrm{i}k\phi}\langle\tilde{a}^{k}\rangle+\mathrm{H.c.}\,,$ (6)
@@
where the operators $\tilde{a}^{k}=\sum_{n=0}^{\infty}\outerproduct{n}{n+k}$ | capture information about the coherence generation and phase localization. To resolve the phase information of the two limit cycles individually, we define truncated operators $\tilde{a}_{\alpha}$ with $\alpha\in\{\mathrm{in,out}\}$,
$\tilde{a}_{\text{in}}$ $=\sum_{n=0}^{n_{c}-1}\outerproduct{n}{n+1},~{}~{}\tilde{a}_{\text{out}}=\sum_{n=n_{c}}^{\infty}\outerproduct{n}{n+1}\,,$ (7)
@@
as an approximation of operators that act on the respective subspace of each ring. The cutoff Fock number $n_{c}$ is chosen to be the integer closest to $r_{c}^{2}$. We use these $\tilde{a}_{\alpha}$ to define the phase distributions $P_{1}^{\text{in}}$ of the inner and $P_{1}^{\text{out}}$ of the outer ring of a TLC,
@@
$\mathcal{I}^{\text{in}}$ $=\sum_{n=0}^{n_{c}}\outerproduct{n}{n},~{}~{}\mathcal{I}^{\text{out}}=\sum_{n=n_{c}+1}^{\infty}\outerproduct{n}{n}\,,$ (9)
@@
We use the measure $P_{1}^{\alpha}$ to characterize the synchronization properties of the two limit cycles. In Figs. 2(b) and 2(c), the locking phase angle of a driven TLC is shown for fixed drive strength $\Omega=0.25\gamma_{1}$ and varying detuning $\Delta$ and Kerr nonlinearity $K$. Remarkably, for nonzero $\Delta$ and $K$, each limit cycle locks to a distinct phase. Notably, the inner limit cycle responds more strongly to the external drive compared to the outer limit cycle. This deviates from the mean-field result Eq. 3, which predicts a stronger phase sensitivity of the outer limit cycle, consistent with standard single-limit cycle oscillators with $\gamma_{3}=\gamma_{4}=0$ [45]. The higher-order gain and damping channels lead to a distinct dynamical regime that is unique to TLCs.
@@
The phase distribution of two oscillators is obtained by projecting the density matrix onto the tensor products of phase states |$\ket{\phi_{A},\phi_{B}}$ and is defined as
$P_{2}(\phi_{BA})$ $=\int\limits_{0}^{2\pi}\mathrm{d}\phi\,\bra{\phi,\phi_{BA}+\phi}\rho\ket{\phi,\phi_{BA}+\phi}-\frac{1}{2\pi}$
@@
where $\tilde{a}_{j,\alpha}$ and $\mathcal{I}^{\alpha}_{j}$ act on the $j$th oscillator. The measure above allows us to investigate the synchronization between the limit cycles of both oscillators since $(\alpha,\beta)$ can take various combinations: $(\mathrm{in,in})$, $(\mathrm{in,out})$, $(\mathrm{out,in})$, and $(\mathrm{out,out})$.
We now set the dissipation rates $\gamma^{A}_{j}=\gamma^{B}_{j}$ equal, such that the radii of the inner and outer limit cycles of both TLCs are identical. In Figs. 3(a–c), we plot the maxima of the combined synchronization measure $P_{2}^{\alpha,\beta}$. For $\delta=0$ and $K=0$, both oscillators have the same frequencies, and hence, we expect a maximum amount of synchronization. Surprisingly, the synchronization measure for $(\alpha,\beta)=(\mathrm{in,in})$ and $(\alpha,\beta)=(\mathrm{out,out})$ shown in Figs. 3(a) and 3(c) is highly suppressed around $K=0$. This is a signature of the synchronization blockade, where the contribution from the first-order locking vanishes ($\langle\tilde{a}^{\dagger}_{A,\alpha}\tilde{a}^{\phantom{\dagger}}_{B,\alpha}\rangle=0$) due to the cancellation of coherences. Only second-order phase locking can be observed, as indicated by the two maxima in Fig. 3(e). This blockade can also be understood using the mean-field equations of the relative phase,
@@
where the coupling term vanishes for limit cycles with equal radii. The limit cycles with different radii exhibit an Arnold tongue, see Fig. 3(b), signifying synchronization between the $(\mathrm{in,out})$ limit cycles. Thus, both synchronization and blockade effects occur simultaneously in the coupled identical TLC oscillators.
The synchronization blockade is lifted for $K\neq 0$ [22], as shown in Figs. 3(a) and 3(c). For $\delta=0$ and $K>0$ ($K<0$) the relative phase of both oscillators locks to $\phi_{BA}=\pi$ ($\phi_{BA}=0$), see Fig. 3(f). To understand this behavior, we investigate the mean-field equations (see [45] for more details). We examine the phase-locking behavior by expanding Eq. 13 about the radii $r_{1}$ and $r_{2}$ of the limit cycles in the mean-field approximation. There are two cases: (i) equal radii $r_{j}=r_{x}+r_{j,1}g/\gamma_{1}$ where $j\in\{A,B\}$, $x,y\in\{1,2\}$, and $x\neq y$,
@@
For $\delta=K=0$, these mean-field equations lead to (i) bistable locking at $\phi_{BA}=0,\pi$ and (ii) locking to a single phase value, like in the full quantum system shown in Figs. 3(d) and 3(e). In contrast to the quantum case, the relative phase in Eq. 15 locks to $\phi_{BA}=\pi/2$ for $(\alpha,\beta)=(\mathrm{in},\mathrm{out})$ and to $\phi_{BA}=-\pi/2$ for $(\alpha,\beta)=(\mathrm{out},\mathrm{in})$. For $\delta=0$ and $K>0$ ($K<0$), the equation of motion Eq. 14 of the relative phase of equal limit cycles has a single stable solution at $\phi_{BA}=\pi$ ($\phi_{BA}=0$), see Fig. 3(f). For different limit cycles, however, $K>0$ ($K<0$) in Eq. 15 leads to a shift of the locking phase towards $\phi_{BA}=0$ ($\phi_{BA}=\pi$). The mean-field analysis is suitable to predict the locking of the relative phases of (i) equal limit cycles of two coupled TLCs but fails to describe the locking of (ii) different limit cycles. An explanation might be that in the quantum Liénard system, both limit cycles are not strictly separated like the basins of attraction in the classical analogue. Therefore, locking mechanisms of different pairs of limit cycles of two TLC oscillators interplay.
For our choice of parameters $n_{c}=2$ and $\tilde{a}_{\mathrm{out}}\approx\tilde{a}$, $P_{2}^{\mathrm{out},\mathrm{out}}$ shown in Fig. 3(c) behaves qualitatively similar to the standard synchronization measure $P_{2}$ defined in Eq. 11. The existence of the blockade at $K=0$ is therefore also confirmed by $P_{2}$, see Fig. 3(e). In [45], we furthermore explore the mutual information as a measure of synchronization [17] and find its behavior to be qualitatively similar to our choice of synchronization measure $P_{2}^{\alpha,\beta}$. However, the blockade at $K=0$ is not well captured.
Experimental realization— The effects discussed here can be potentially observed in a trapped-ion experiment similar to [50] that demonstrated quantum synchronization of a phonon laser to an external signal. The setup consists of a calcium and a beryllium ion in a radio-frequency trap that share a common harmonic mode of motion which is denoted by the annihilation operator $a$. To realize $n$th-order gain ($m$th-order damping), one has to implement a sideband heating (cooling) laser that is detuned from a particular transition in the ion energy-level scheme [51]. If this detuning equals $n$ ($-m$) times the energy of the harmonic mode and assuming fast ion decay with respect to the time scales of the motion in the trap, an effective jump operator $L=a^{\dagger n}$ ($L=a^{m}$) is realized. For each dissipator in Eq. 1, a distinct ion transition has to be chosen. Therefore, to realize the four gain and damping channels of a TLC, two spin transitions per ion have to be driven with one of the four red and blue sideband lasers each.
@@
Our setup can be extended by incorporating higher-order gain and damping channels, leading to multiple local minima in the effective potential and, therefore, multiple limit cycles. Another choice of dissipation channels that are more localized in Fock space has been studied in [52]. Employing such channels will lead to multiple effective few-level limit cycles in Fock space centered at various Fock numbers. Future directions also include the study of minimal examples, e.g., spin-2 oscillators where both |±$\ket{\pm 1}$ are stabilized, as well as networks of TLCs. The analysis of multiple limit cycles within a single steady state opens a promising avenue within the field of quantum synchronization with potential applications in quantum sensing and entanglement generation.

### v2

@@ Abstract
Introduction—Synchronization phenomena of self-sustained oscillating systems have been studied for centuries [1, 2] and remain an active research topic of nonlinear dynamics [3, 4, 5, 6]. Self-sustained oscillators exhibit limit cycles, i.e., closed phase-space trajectories that the system converges to regardless of its initial state. Such limit-cycle oscillators can align quantities like their phase or frequency with other oscillators as well as with external signals. Quantum analogues of limit-cycle oscillators have been studied in various settings ranging from quantum van der Pol oscillators [7, 8, 9, 10, 11, 12, 13] to few-level quantum oscillators [14, 15, 16, 17]. While some features of quantum synchronization are inherited from classical synchronization, others originate from genuine quantum properties like entanglement [18, 19, 20, 21, 22, 23] and (destructive) interference that can lead to a synchronization blockade [24, 25, 26]. Understanding quantum synchronization has implications for quantum sensing [27], quantum thermodynamics [28, 29, 30], quantum networks [31], and time crystals [32, 33, 34]. First steps toward its experimental realization have been taken on several platforms including cold atoms [35], nuclear spins [36], trapped ions [37], and superconducting circuits [38].
@@
While there has been an increasing interest in quantum systems featuring multiple separate limit cycles [42, 43, 44, 45, 46, 47], their synchronization properties have not been studied. In this work, we introduce a Liénard system where two limit cycles coexist in a single steady state regardless of the initial conditions and investigate their synchronization behavior. We call this system a twin limit cycle (TLC): it is characterized by a double ring-like structure in phase space as sketched in Fig. 1(b). The location of the minima (maxima) of the effective potential of the corresponding classical Liénard system is indicated by the dashed (dotted) rings. Our setup can be extended to host multiple limit cycles. We examine the synchronization of a single TLC under an external coherent drive and find that the limit cycles exhibit different locking behaviors. Furthermore, for two coupled identical TLCs, both synchronization and blockade effects coexist, an apparent paradoxical interplay unattainable with standard limit cycles. To distinguish the contributions of individual limit cycles of a TLC, we define new finer measures of quantum synchronization. Finally, we outline an experimental setup to implement our model.
Figure 1: Illustration of a system with multiple limit cycles. (a) Sketch of the effective potential $V(r)$ of a classical Liénard system featuring two basins of attraction separated at $r_{c}$. (b) Wigner function of a twin limit cycle, i.e., a quantum Liénard system, with $\langle a\rangle=r\mathrm{e}^{\mathrm{i}\phi}$. The dashed (dotted) rings with radii $r_{1}$ and $r_{2}$ ($r_{c}$) correspond to the local minima (maximum) of the effective potential obtained from the mean-field equations of motion, see Eq. 4.
Model— We consider a coherently driven anharmonic quantum oscillator subject to incoherent first and third-order pumping, along with second and fourth-order damping. These dissipative processes stabilize two concentric limit cycles, see Fig. 1(b). The dynamics in the rotating frame of the drive is described by the master equation
@@
where $\mathcal{D}[L]\rho=L\rho L^{\dagger}-(L^{\dagger}L\rho+\rho L^{\dagger}L)/2$ is the Lindblad dissipator, $H_{0}=\Delta a^{\dagger}a^{\phantom{\dagger}}+Ka^{\dagger 2}a^{2}$, $H_{d}=\Omega(a+a^{\dagger})$, and $a$ ($a^{\dagger}$) denote the annihilation (creation) operators of the oscillator. The detuning between the TLC and the drive is denoted by $\Delta$, $K$ parametrizes the Kerr nonlinearity, and $\Omega$ denotes the strength of the drive. The rates $\gamma_{j}$ correspond to incoherent gain (odd $j$) and damping (even $j$). The model simplifies to the paradigmatic quantum van der Pol oscillator with a single limit cycle for $\gamma_{3}=\gamma_{4}=0$. Additional incoherent processes of higher order, $\mathcal{D}[a^{\dagger n}]\rho$ and $\mathcal{D}[a^{m}]\rho$, with $n(m)$ being odd(even), can be included to obtain multiple limit cycles.
We begin by examining the semiclassical limit to obtain an approximation to the steady state of the quantum system. The mean-field equations of the effective classical Liénard model can be derived by performing a cumulant expansion to first order. Setting $\langle a\rangle=r\mathrm{e}^{\mathrm{i}\phi}$ results in
@@
$\dot{\phi}=$ $-\Delta-2Kr^{2}-\frac{\Omega}{r}\cos(\phi)\,.$ (3)
Since we are interested in the case of two stable limit cycles, we choose $\gamma_{j}$ such that the right-hand side of Eq. 2 exhibits three real zeros $r_{1}<r_{c}<r_{2}$ at vanishing drive $\Omega=0$,
$\dot{r}=$ $r(r_{1}^{2}-r^{2})(r_{c}^{2}-r^{2})(r_{2}^{2}-r^{2})2\gamma_{4}\equiv-\partial_{r}V(r)\,.$ (4)
Here, $r_{1}$ and $r_{2}$ are the stable solutions, and $r_{c}$ is the unstable solution of the mean-field equations, and $V(r)$ is the effective potential, see Fig. 1(a). In a classical system, $r_{c}$ separates the two basins of attraction. An initial state with $r<r_{c}$ ($r>r_{c}$) will therefore converge to $r_{1}$ ($r_{2}$).
@@
In a classical Liénard system of two limit cycles, both basins of attraction are separated at $r_{c}$. However, if extrinsic noise is added, trajectories can cross this boundary. In the language of the effective potential shown in Fig. 1(a), noise-induced jumps in $r$ have to overcome the potential barriers $\Delta V_{j}$ to “tunnel” between both limit cycles. In our quantum setup, the steady state is a combination of two distinct quantum limit cycles. Considering quantum trajectories, the system tunnels between the two limit cycles due to inherent quantum noise, see Supplemental Material [48] for a detailed discussion.
@@
Figure 2: Drive-induced phase locking of a twin limit cycle (TLC). (a) Wigner function for $\Omega=8\gamma_{1}$ and $\Delta=K=0$ showing phase localization of both limit cycles, i.e., a maximum at $\phi=\mathrm{arg}(\langle a\rangle)=\mathrm{arg}(\Omega)-\pi/2=-\pi/2$. Here, dashed (dotted) rings correspond to stable (unstable) solutions of the mean-field equations of the undriven limit cycles, see Eq. 4. The solid orange curves are contour lines at $0.0225$. Note the power-law color scale. In panels (b) and (c), solid curves denote the maximum of $P_{1}^{\alpha}$ with $\alpha\in\{\mathrm{in,out}\}$, and dashed curves denote $\text{arg}(\langle\tilde{a}_{\alpha}\rangle)$. Varying $\Delta$ and $K$, the inner limit cycle exhibits a larger phase shift than the outer one. This is opposite to the behavior of a standard quantum van der Pol oscillator, see Supplemental Material [48]. Here, $\Omega=0.25\gamma_{1}$ for both (b) $K=0$ and (c) $\Delta=0$. The dissipation rates for (a)–(c) are $\gamma_{2}=2.5\gamma_{1}$, $\gamma_{3}=1.04\gamma_{1}$, and $\gamma_{4}=0.096\gamma_{1}$.
[…]
$P_{1}(\phi)$ $=\frac{1}{2\pi}\sum_{k=1}^{\infty}\mathrm{e}^{-\mathrm{i}k\phi}\langle\tilde{a}^{k}\rangle+\mathrm{H.c.}\,$, (6)
@@
where the operator powers $|\tilde{a}^{k}=\sum_{n=0}^{\infty}\outerproduct{n}{n+k}$ capture information about the coherence generation and phase localization. To resolve the phase information of the two limit cycles individually, we define truncated operators $\tilde{a}_{\alpha}$ with $\alpha\in\{\mathrm{in,out}\}$,
[…]
@@
as an approximation of operators that act on the respective subspace of each ring. The cutoff Fock number $n_{c}$ is chosen to be the integer closest to $r_{c}^{2}$. We use powers of these $\tilde{a}_{\alpha}$ to define the phase distributions $P_{1}^{\text{in}}$ of the inner and $P_{1}^{\text{out}}$ of the outer ring of a TLC,
@@
[…]
@@
We use the measure $P_{1}^{\alpha}$ to characterize the synchronization properties of the two limit cycles. In Figs. 2(b) and 2(c), the locking phase angle of a driven TLC is shown for fixed drive strength $\Omega=0.25\gamma_{1}$, varying detuning $\Delta$, and Kerr nonlinearity $K$. Remarkably, for nonzero $\Delta$ and $K$, each limit cycle locks to a distinct phase. Notably, the inner limit cycle responds more strongly to the external drive compared to the outer limit cycle. This deviates from the mean-field result Eq. 3, which predicts a stronger phase sensitivity of the outer limit cycle, consistent with standard single-limit-cycle oscillators with $\gamma_{3}=\gamma_{4}=0$ [48]. The higher-order gain and damping channels lead to a distinct dynamical regime that is unique to TLCs. The overlap of the inner and outer limit cycles of a TLC leads to tunneling and leakage of information between them. This is a qualitatively different behavior compared with the classical analogue following the mean-field equations. In addition to the locking of the two limit cycles of a driven TLC to distinct phases shown in Fig. 2, we discuss their frequency synchronization in the Supplemental Material [48].
@@
[…]
@@
where $\tilde{a}_{j,\alpha}$ and $\mathcal{I}^{\alpha}_{j}$ are the truncated operators $\tilde{a}_{\alpha}$ and the unit operators $\mathcal{I}^{\alpha}$ that act on the $j$th oscillator. The measure above allows us to investigate the synchronization between the limit cycles of both oscillators since $(\alpha,\beta)$ can take various combinations: $(\mathrm{in,in})$, $(\mathrm{in,out})$, $(\mathrm{out,in})$, and $(\mathrm{out,out})$.
We now set the dissipation rates $\gamma^{A}_{j}=\gamma^{B}_{j}$ equal, such that the radii of the inner and outer limit cycles of both TLCs are identical. In Figs. 3(a) to 3(c), we plot the maxima of the combined synchronization measure $P_{2}^{\alpha,\beta}$. For $\delta=0$ and $K=0$, both oscillators have the same frequencies, and hence, we expect a maximum amount of synchronization. Surprisingly, the synchronization measure for $(\alpha,\beta)=(\mathrm{in,in})$ and $(\alpha,\beta)=(\mathrm{out,out})$ shown in Figs. 3(a) and 3(c) is highly suppressed around $K=0$. This is a signature of the synchronization blockade, where the contribution from the first-order locking vanishes ($\langle\tilde{a}^{\dagger}_{A,\alpha}\tilde{a}^{\phantom{\dagger}}_{B,\alpha}\rangle=0$) due to the cancellation of coherences. Only second-order phase locking can be observed, as indicated by the two maxima in Fig. 3(e). This blockade can also be understood using the mean-field equations of the relative phase,
@@
where the coupling term vanishes for limit cycles with equal radii. The limit cycles with different radii exhibit an Arnold tongue, see Fig. 3(b), signifying synchronization between the $(\mathrm{in,out})$ limit cycles. Thus, both synchronization and blockade effects occur simultaneously in the coupled identical TLC oscillators, a behavior not known in classical analogues in the absence of noise.
The synchronization blockade is lifted for $K\neq 0$ [24], as shown in Figs. 3(a) and 3(c). For $\delta=0$ and $K>0$ ($K<0$) the relative phase of both oscillators locks to $\phi_{BA}=\pi$ ($\phi_{BA}=0$), see Fig. 3(f). To understand this behavior, we investigate the mean-field equations (see Supplemental Material [48] for more details). We examine the phase-locking behavior by expanding Eq. 13 about the radii $r_{1}$ and $r_{2}$ of the limit cycles in the mean-field approximation. There are two cases: (i) equal radii $r_{j}=r_{x}+r_{j,1}g/\gamma_{1}$ where $j\in\{A,B\}$, $x,y\in\{1,2\}$, and $x\neq y$,
@@
For $\delta=K=0$, these mean-field equations lead to (i) bistable locking at $\phi_{BA}=0,\pi$ and (ii) locking to a single phase value, like in the full quantum system shown in Figs. 3(d) and 3(e). In contrast to the quantum case, the relative phase in Eq. 15 locks to $\phi_{BA}=\pi/2$ for $(\alpha,\beta)=(\mathrm{in},\mathrm{out})$ and to $\phi_{BA}=-\pi/2$ for $(\alpha,\beta)=(\mathrm{out},\mathrm{in})$. For $\delta=0$ and $K>0$ ($K<0$), the equation of motion Eq. 14 of the relative phase of equal limit cycles has a single stable solution at $\phi_{BA}=\pi$ ($\phi_{BA}=0$), see Fig. 3(f). For different limit cycles, however, $K>0$ ($K<0$) in Eq. 15 leads to a shift of the locking phase toward $\phi_{BA}=0$ ($\phi_{BA}=\pi$). The mean-field analysis is suitable to predict the locking of the relative phases of (i) equal limit cycles of two coupled TLCs but fails to describe the locking of (ii) different limit cycles. An explanation might be that in the quantum Liénard system, both limit cycles are not strictly separated like the basins of attraction in the classical analogue. Therefore, locking mechanisms of different pairs of limit cycles of two TLC oscillators interplay.
For our choice of parameters $n_{c}=2$ and $\tilde{a}_{\mathrm{out}}\approx\tilde{a}$, $P_{2}^{\mathrm{out},\mathrm{out}}$ shown in Fig. 3(c) behaves qualitatively similar to the standard synchronization measure $P_{2}$ defined in Eq. 11. The existence of the blockade at $K=0$ is therefore also confirmed by $P_{2}$, see Fig. 3(e). In the Supplemental Material [48], we furthermore explore the mutual information as a measure of synchronization [19] and find its behavior to be qualitatively similar to our choice of synchronization measure $P_{2}^{\alpha,\beta}$. However, the blockade at $K=0$ is not well captured.
Experimental realization— The effects discussed here can be potentially observed in a trapped-ion experiment similar to [53] that demonstrated quantum synchronization of a phonon laser to an external signal. The setup consists of a calcium and a beryllium ion in a radio-frequency trap that share a common harmonic mode of motion which is denoted by the annihilation operator $a$. To realize $n$th-order gain ($m$th-order damping), one has to implement a sideband heating (cooling) laser that is detuned from a particular transition in the ion energy-level scheme [54]. If this detuning equals $n$ (m-m) times the energy of the harmonic mode and assuming fast ion decay with respect to the timescales of the motion in the trap, an effective jump operator $L=a^{\dagger n}$ ($L=a^{m}$) is realized. For each dissipator in Eq. 1, a distinct ion transition has to be chosen. Therefore, to realize the four gain and damping channels of a TLC, two spin transitions per ion have to be driven with one of the four red and blue sideband lasers each.
@@
[…]